OpenStudy (anonymous):

If a 375 mL sample of water was cooled from 37.5 C to 0 C how much heat wad lost in joules

OpenStudy (michele_laino):

please, you have to apply the formula below: \[Q=c*m*\Delta t\] where Q is the energy lost, c is the specific heat of water, m is the mass of water involved, so m=3.75 *10^-1 Kg c=4,184 J/(Kg*°C) delta t=37.5 °C

OpenStudy (anonymous):

Taking density of water as 1000kg/m3. Mass of water would be 0.375kg. So, heat lost would be \[H=mC DeltaT\] H=0.375*4184*37.5 = 58837.5J

OpenStudy (michele_laino):

please, you have to perform the calculation using the above data